.1.What will my investment of R100 500 be worth if my bank pays me simple interest of 9% per annum and I invest my money for 180 days?
1
Expert's answer
2020-11-25T12:18:29-0500

Your investment worth after 180 days:

= Investment + Interest

"= 100500 + (100500 \times 0.09 \times \frac{180}{365})"

= 100500 + 4461

= 100961

Answer: R100961